my question is
at what annual rate compounded continuously will $1,000 have to be invested to amount to $2,500 in 10 years?
i know the answer is 9.16% an you use the formula A=Pe^rt
but how do you find the answer on the caculator and what numbers go where?

Answer by nerdybill(6962)

at what annual rate compounded continuously will $1,000 have to be invested to amount to $2,500 in 10 years?
i know the answer is 9.16% an you use the formula A=Pe^rt
but how do you find the answer on the caculator and what numbers go where?
.
Since:
A=Pe^(rt)
Plug in what you know:
2500=1000e^(r*10)
.
Solving for r:
2500=1000e^(r*10)
2500/1000 = e^(r*10)
ln(2500/1000) = r*10
ln(2500/1000)/10 = r
.
So on your calculator:
Hit the "Ln" button
Type in "2500/1000"
Hit the ")"
Hit the "divide" key
Type in "10"
Hit the "Enter" key
.
r = .0916 or 9.16%